OpenStudy (anonymous):

graph V(t)=Cos(t^2-2) in the interval of 0 to pi -find average velocity -find the point of maximum velocity and point of minimum velocity (Plz help i dont know how to do this problem and would appreciate an explanation so i know how to do problems in the future similar to this)

OpenStudy (anonymous):

Are you in calculus? Have you done integrals?

OpenStudy (anonymous):

\[\frac{ 1 }{ \pi-0 } \int\limits_{0}^{\pi} f(x) dx\] = average of f(x) over x=0..pi